# Forecasting Trends in Food Security: a Reservoir Computing Approach

### v2

@@ Discussion
This manuscript applies several purely quantitative time series prediction methodologies to the problem of forecasting insufficient food consumption, based on the data collected by the World Food Programme’s real-time monitoring system. In direct comparison with LSTM, CNN and ARIMA we find Reservoir Computing to be the most suitable for this task in terms of performance, resource requirements, robustness and ease of use. We further proved this approach to surpass a previous attempt using XGBoost[21]. The Reservoir Computing algorithm has demonstrated to be a solid framework for the construction of a purely quantitative early warning system, owing to several advantageous features. Its ability to process multidimensional time series and learn complex nonlinear relations in target and features gives it the advantage over simple statistical models like ARIMA. The iterative prediction scheme allows the use of exogenous variables like seasonal and religious calendars where future values are already known as well as make implicit forecasts for those where this is not the case. The relative simplicity of the approach mitigates the danger of overfitting when compared to the more sophisticated Deep Learning methodologies CNN and LSTM, in line with what was shown in [47]. The limited resources needed for the training process allow an ensemble approach, the size of which can be fined tuned and adjusted based on the available computational resources and accuracy requirements of the field and context of application. Furthermore, we saw how the RC methodology was able to accommodate a degree of class imbalance in the training data. Indeed the above benchmark performance shown by the RC model does not depend on the majority of stable curves that form the training set, but is actually more evident in the subset of curves that represent dramatic increases in the levels of insufficient food consumption. These findings suggest the potential advantages of RC over models of this kind in data-scarce contexts, even when confronted with high-dimensional, noisy, and authentic datasets. The ensemble architecture appears to perform well without requiring specific modifications for large input dimensions, as seen in approaches like Local States [52].
@@
Future improvements of the forecasting methodology will concentrate on four main pillars: extending the forecasting horizon, tailoring the methodology to detect severe deterioration’s of food security levels, developing methods to explain the role of the input variables and implementing the methodology in a production environment. Initial results obtained on 90- day forecasts confirm the RC algorithm as the preferred choice with respect to other benchmarks, despite a general increase of the errors (see Supplementary Material). Aggregating the underlying data to weekly or monthly time granularity, reducing the number of data points that need to be forecasted and smoothing uninformative fluctuations, seems like a promising possibility. Increasing the target’s time granularity would also eliminate the need for interpolation and other pre-processing steps thereby facilitating the deployment of the approach. Introducing a custom weighting system in the training process and duplicating data splits which show severe deterioration will be attempted to focus the methodology on those curves that are more important to anticipate. Also testing other and more modern architectures of Reservoir Computing like the Next Generation RC[53] or the Trend-Seasonality decomposition RC [47] variation might represent a valid choice to improve overall performances and extend the horizon. Finally, allowing the algorithms to autonomously select the most informative subset of features for each forecast, as done in this manuscript, marks an initial effort to comprehend the underlying drivers influencing the predictions. A comprehensive analysis of the results and the continued development of this approach will be crucial for enhancing the explanatory power of the methodology.
@@
In conclusion, this research lays the groundwork for a dedicated quantitative tool for modeling and forecasting food insecurity. Our approach complements existing methodologies with its year-round versatility by offering real time updates on future levels of food insecurity. By leveraging the advantageous features of real-time monitoring, the methodology proves particularly valuable for foreseeing and quantifying the impact of events that occur between the releases of other monitoring systems. Our work presents the potential to furnish humanitarian organizations with the capability to rapidly institute new early warning systems to preemptively help vulnerable communities. The utilization of extant and expanding datasets in conjunction with the low requirement of computational power renders the proposed approach simple and economical to implement. As such, it could offer much-needed technical support to humanitarian actors in their efforts to anticipate, quantify and promptly address impending food crises.
@@ Methods
Data
Figure 4: Feature groupings that were considered in the grid search for the RC, CNN and LSTM models to implement a dynamic feature selection.(a) Schema of how the features were classified into 5 different feature groupings. (b-d) frequency with which each feature grouping was selected for in (a) the RC model, (b) the LSTM model and (c) the CNN model.
@@
Table 1: Variables used in the model for each of the four pilot countries. Time resolution of the data is indicated under frequency. The level of spatial aggregation is shown under spatial resolution. The last row shows the total number of input variables used in each model, which is summed over all sub-national time series available for the specified features.
Rainfall and Vegetation: Five rainfall and vegetation features were constructed, sourced from WFP’s Seasonal Explorer [61] and are originally provided by CHIRPS [62] (rainfall) and MODIS[63] (vegetation) satellite imagery. These include the amount of rainfall in mm, the 1- and 3-month anomaly of rainfall compared to historical seasonal behavior. Similarly, the NDVI value and its 1-month anomaly are used. The NDVI (Normalized Difference Vegetation Index) is a measure for the amount of vegetation in a region. From these WFP derives regional seasonal calendars which mark agriculturally relevant seasons. They can be considered a binary time series of either 0 or 1. All of these variables are given in 10-day intervals ("dekads") for every region.
Conflict: To represent conflict as a driver of food insecurity we use data from the Armed Conflict Location & Event Data Project (ACLED) [64] which is a publicly available repository providing real-time and historical data on political violence and protest events in nearly 100 countries [65]. We build our features from the fatalities reported for events in an administrative region of the categories Battle, Violence against Civilians and Explosions/Remote Violence by summing them over the last 90 days in order to create daily time series.
$ALPS=({Price}_{t}-\hat{Price}_{t})/\sigma_{\epsilon}$ (3)
Figure 5: Prevalence of people with insufficient food consumption averaged over all available regions of Yemen in blue. Time periods of Ramadan shown by areas shaded red.
Training Procedure
Results directly derived from the grid search are shown in figure 6. The importance of the spectral radius and the input strength in RC is confirmed in agreement with the literature. For CNN we can see that differencing improves the the performance significantly, while other hyperparameters like e.g. the number of filters only have a small impact. For LSTM we identify the learning rate, the number of units, the dropout rate and the length of lookback as having a significant impact. We can see that it is generally impossible to define a single hyperparameter as good or bad since they are usually interdependent. Nevertheless, the fact that high dropout rates are favored in some cases in LSTM indicates a tendency towards overfitting. Overall, it is observed that RC stands out as the model for which it is easiest to identify generally reasonable hyperparameter configurations that yield satisfactory performance across all splits. This implies that satisfactory forecasts could even be made with less frequent retuning of the model’s hyperparameters.
  1. 1.
  2. 2.
    Training the Model: The model is trained on the training data.
  3. 3.
  4. 4.
Figure 6: Additional results on the influence of different hyperparameters derived from the grid search on Yemen. The value shown is always the median RMSE over all splits and unspecified hyperparameters. (a-b): RC results for spectral radius and input strength, which are often considered the most important hyperparameters in RC. Results are separated into runs with and without differencing to showcase varying behavior. (d-f): CNN results on differencing, learning rate and number of filters. (g-i): LSTM results on learning rate, number of units, dropout rate and length of lookback.
@@
Table 2: Ranges of RC hyperparameters to be searched during tuning procedure.
@@
Table 3: Ranges of LSTM hyperparameters to be searched during tuning procedure.
@@
Table 4: Ranges of CNN hyperparameters to be searched during tuning procedure.
@@
Table 5: Ranges of ARIMA hyperparameters to be searched during hyperparameter tuning.
Data Availability
The data that support the findings of this study are available from the corresponding author, D.P.
Code Availability
The code that supports the findings of this study is available from the corresponding author, D.P.
Acknowledgements
We would like to thank Kyriacos Koupparis and Arif Husain for the continuous support. We are also grateful to Elisa Omodei and Pietro Foini for sharing their data and their knowledge on the XGBoost approach. We also want to express our gratitude to Will Fein from Microsoft AI for Good for his suggestions that improved our training procedure. This work was supported by the World Food Programme Innovation Accelerator.
Author contributions
C.R. conceived and supervised the research. J.H. and D.P. designed the study. J.H. and D.P. conducted the calculations. All authors interpreted and evaluated the findings. All authors wrote and reviewed the manuscript.
Competing interests
All authors declare no competing interests.
Correspondence

### v3

@@ Discussion
This manuscript applies several purely quantitative time series prediction methodologies to the problem of forecasting insufficient food consumption of a horizone of 60 days, based on the data collected by the World Food Programme’s real-time monitoring system. In direct comparison with LSTM, CNN and ARIMA we find Reservoir Computing to be the most suitable for this task in terms of performance, resource requirements, robustness and ease of use. We further proved this approach to surpass a previous attempt using XGBoost[32] (See supplementary material section 6).
@@
The Reservoir Computing algorithm has demonstrated to be a solid framework for the construction of a purely quantitative early warning system, owing to several advantageous features. Its ability to process multidimensional time series and learn complex nonlinear relations in target and features gives it the advantage over simple statistical models like ARIMA. The iterative prediction scheme allows the use of exogenous variables like seasonal and religious calendars where future values are already known as well as make implicit forecasts for those where this is not the case. The relative simplicity of the approach mitigates the danger of overfitting when compared to the more sophisticated Deep Learning methodologies CNN and LSTM, in line with what was shown in [44]. The limited resources needed for the training process allow an ensemble approach, the size of which can be fined tuned and adjusted based on the available computational resources and accuracy requirements of the field and context of application. Furthermore, we saw how the RC methodology was able to accommodate a degree of class imbalance in the training data. Indeed the above benchmark performance shown by the RC model does not depend on the majority of stable curves that form the training set, but is actually more evident in the subset of curves that represent dramatic increases in the levels of insufficient food consumption. These findings suggest the potential advantages of RC over models of this kind in data-scarce contexts, even when confronted with high-dimensional, noisy, and authentic datasets. The ensemble architecture appears to perform well without requiring specific modifications for large input dimensions, as seen in approaches like Local States [49]. Although the comparisons with methodologies based on different target variables is not straightforward, its worth noting that the performances measured with RMSE are in line or better of what presented in other works. In [26] with IPC3+ target the RMSE is of $13\%$, while in [50], the best model performs between $5\%$ and $13\%$ RMSE, although on different time horizons. In [25] the performances vary considerably for different countries up too a maximum of $14.7\%$ RMSE for near term forecasts and to $17.5\%$ on medium term. Other works used different metrics to measure the errors, such as $R^{2}$, making it more difficult, to compare with what obtained in this work.
@@
The results shown in this manuscript indicate that a forecasting methodology based on the RC algorithm is able to predict 60 days into the future the levels of insufficient food consumption with an average RMSE of 5 percentage points. This increases to around 10 percentage points when only focusing on the subset of curves that show a severe deterioration. From the classification exercise we see that the RC correctly forecasts the type of behaviour once every two forecasts (Total accuracy 0.48), and the number goes down to once every three forecasts when only considering classifying future behaviour as deteriorating of not (accuracy 0.33), or improving or not (accuracy 0.36), according to our definitions. We believe that this is solid starting point for the development of an early warning system based on the RC forecasting algorithm. More work will be done both to improve the general performances and to tailor the methodology to on deteriorating curves. To achieve this goal we will concentrate our future work on four main pillars: extending the forecasting horizon, tailoring the methodology to detect severe deterioration’s of food security levels, developing methods to explain the role of the input variables and implementing the methodology in a production environment. Initial results obtained on 90- day forecasts confirm the RC algorithm as the preferred choice with respect to other benchmarks, despite a general increase of the errors (see Supplementary Material Section 1). Aggregating the underlying data to weekly or monthly time granularity, reducing the number of data points that need to be forecasted and smoothing uninformative fluctuations, seems like a promising possibility. This would also eliminate the need for interpolation and other pre-processing steps on secondary variables, thereby facilitating the deployment and maintanance of the approach. Introducing a custom weighting system in the training process and duplicating data splits which show severe deterioration will be attempted to focus the methodology on deteriorating curves, that are the most important to anticipate. Also testing other and more modern architectures of Reservoir Computing like the Next Generation RC[51] or the Trend-Seasonality decomposition RC [44] variation might represent a valid choice to improve overall performances and extend the horizon. In section 2 of the supplementary material we have described a procedure to compute the confidence intervals of the forecasted curves. The confidence intervals may be added considering ±2-standard errors confidence bands. This methodology can be applied to all algorithms including RC. Finally, allowing the algorithms to autonomously select the most informative subset of features for each forecast, as done in this manuscript, marks an initial effort to comprehend the underlying drivers influencing the predictions. A comprehensive analysis of the results and the continued development of this approach will be crucial for enhancing the explanatory power of the methodology.
@@
In conclusion, this research lays the groundwork for a dedicated quantitative tool for modeling and forecasting food insecurity. Our approach complements existing methodologies with its year-round versatility by offering real time updates on future levels of food insecurity. By leveraging the advantageous features of real-time monitoring, the methodology proves particularly valuable for foreseeing and quantifying the impact of events that occur between the releases of other monitoring systems. Our work presents the potential to furnish humanitarian organizations with the capability to rapidly institute early warning systems to preemptively help vulnerable communities. The utilization of extant and expanding datasets in conjunction with the low requirement of computational power renders the proposed approach simple and economical to implement. As such, it could offer much-needed technical support to humanitarian actors in their efforts to anticipate, quantify and promptly address impending food crises.
@@ Methods
Data
Rainfall and Vegetation: Five rainfall and vegetation features were constructed, sourced from WFP’s Climate Explorer [52] and are originally provided by CHIRPS [53] (rainfall) and MODIS[54] (vegetation) satellite imagery. These include the amount of rainfall in mm and the NDVI (Normalized Difference Vegetation Index) value, which is a measure for the amount of vegetation in a region. In addition the anomalies of both indicators are included. The anomaly is calculated by dividing the average value of a variable during a period (e.g. one month) by the historical average for the same period in all previous years. To inform the model about the comparison of current and historical seasonal behavior we include the 1- and 3-month anomaly of rainfalland the 1-month anomaly of NDVI. From these WFP derives regional seasonal calendars which mark agriculturally relevant seasons. They can be considered a binary time series of either 0 or 1. All of these variables are given in 10-day intervals ("dekads") for every region.
Conflict: To represent conflict as a driver of food insecurity we use data from the Armed Conflict Location & Event Data Project (ACLED) [55] which is a publicly available repository providing real-time and historical data on political violence and protest events in nearly 100 countries [56]. We build three features from the fatalities reported for events in an administrative region of the categories Battle, Violence against Civilians and Explosions/Remote Violence by summing each of them in a rolling window over the last 90 days in order to create daily time series for each subnational region.
$ALPS=({Price}_{t}-\hat{Price}_{t})/\sigma_{\epsilon}$ (1)
@@
Figure 7: Grid Search Results: Additional results on the influence of different hyperparameters derived from the grid search on Yemen. The value shown is always the median RMSE over all splits and unspecified hyperparameters. (a-b): RC results for spectral radius and input strength, which are often considered the most important hyperparameters in RC. Results are separated into runs with and without differencing to showcase varying behavior. (d-f): CNN results on differencing, learning rate and number of filters. (g-i): LSTM results on learning rate, number of units, dropout rate and length of lookback.
Training Procedure
Results directly derived from the grid search are shown in figure 7. The importance of the spectral radius and the input strength in RC is confirmed in agreement with the literature. For CNN we can see that differencing improves the performance, while other hyperparameters like e.g. the number of filters only have a small impact. For LSTM we identify the learning rate, the number of units, the dropout rate and the length of lookback as having a strong impact. We can see that it is generally impossible to define a single hyperparameter as good or bad since they are usually interdependent. Nevertheless, the fact that high dropout rates are favored in some cases in LSTM indicates a tendency towards overfitting. Overall, it is observed that RC stands out as the model for which it is easiest to identify generally reasonable hyperparameter configurations that yield satisfactory performance across all splits. This implies that satisfactory forecasts could even be made with less frequent retuning of the model’s hyperparameters.
  1. 1.
  2. 2.
    Training the Model: The model is trained on the training data.
  3. 3.
  4. 4.
Data Availability
The data used in this study is available in public repository https://github.com/ducciopiovani/FamPredAI
Code Availability
The code used in this study is available in public repository https://github.com/ducciopiovani/FamPredAI
Acknowledgements
We would like to thank Arif Husain for the continuous support and guidance. We are also grateful to Elisa Omodei and Pietro Foini for sharing their data and their knowledge on the XGBoost approach. We also want to express our gratitude to Will Fein from Microsoft AI for Good for his suggestions that improved our training procedure. This work was supported by the World Food Programme Innovation Accelerator.
Author contributions
C.R. conceived and supervised the research, D.P. and J.H. designed the study. J.H., I.L. and D.P. conducted the calculations. J.H., C.R, G.M, A.B, K.K, I.L and D.P. interpreted and evaluated the findings. J.H., C.R, G.M, A.B, K.K, I.L and D.P. wrote and reviewed the manuscript.
Competing interests
All authors declare no competing interests.
Correspondence
@@ Acknowledgements
Table 1: Input Data: Variables used in the model for each of the four pilot countries. Time resolution of the data is indicated under frequency. The level of spatial aggregation is shown under spatial resolution. The last row shows the total number of input variables used in each model, which is summed over all sub-national time series available for the specified features.
@@
Table 2: RC Grid Search Parameters: ranges of RC hyperparameters to be searched during tuning procedure.
@@
Table 3: LSTM Grid Search Parameters: ranges of LSTM hyperparameters to be searched during tuning procedure.
@@
Table 4: CNN Grid Search Parameters: ranges of CNN hyperparameters to be searched during tuning procedure.
@@
Table 5: Arima Grid Search Parameters: ranges of ARIMA hyperparameters to be searched during hyperparameter tuning.